OpenStudy (anonymous):

Use Cramer's rule to determine if the system is inconsistent system or contains dependent equations. 2x + 7 = 8 6x + 3y = 24 A. system is inconsistent B. system contains dependent equations

OpenStudy (apoorvk):

do you know what Cramer's rule is?? Ever heard of it?

OpenStudy (anonymous):

yes i know it but idk how to determine whether its inconsistent or dependent..

OpenStudy (anonymous):

Is it 2x + 7 = 8 or 2x + 7y = 8 for the first equation?

OpenStudy (anonymous):

it says 7 in the question..

OpenStudy (anonymous):

So your equations are 2x + 0y = 1 6x + 3y = 24 x=1/2 3 + 3 y =24 3y = 21 y =7 The system is consistent

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

@eliassaab